Irwin Music Shop gives its customers coupons redeemable for a poster plus a Bo Diddley CD. One coupon is issued for each dollar of sales. On the surrender of 100 coupons and $6.00 cash, the poster and CD are given to the customer. It is estimated that 80% of the coupons will be presented for redemption. Sales for the first period were $700,000, and the coupons redeemed totaled 420,000. Sales for the second period were $840,000, and the coupons redeemed totaled 750,000. Irwin Music Shop bought 20,000 posters at $2.50/poster and 20,000 CDs at $7.50/CD.

Instructions

Prepare the following entries for the two periods, assuming all the coupons expected to be redeemed from the first period were redeemed by the end of the second period.

                        Entry                                                                     Period 1                       Period 2    

  1. To record coupons redeemed

——————————————————————————————————————————

  1. To record estimated liability
Answer #1
Period 1 Period 2
Entry Debit Credit Debit Credit
a. Cash [420000/100 = 4200 x $6]; [750000/100 = 7500 x $6] 25200 45000
Premium liability 11200
Premium Expense [4200 x ($2.50 + $7.50 - $6.00)];[(7500 - 2800) x ($2.50 + $7.50 - $6.00)] 16800 18800
Inventory of Premiums [4200 x ($2.50 + $7.50)];[7500 x ($2.50 + $7.50)] 42000 75000
(To record coupons redeemed)
b. Premium Expense 11200 8080
Premium liability 11200 8080
(To record estimated liability)

Working:

Period 1:
Number of coupons expected to be redeemed = $700000/100 = 7000 x 80% = 5600
Actually redeemed = 4200
Premium liability = [(7000 - 4200) x ($2.50 + $7.50 - $6.00)] = 2800 x $4 = $11200
Period 2:
Number of coupons expected to be redeemed = $840000/100 = 8400 x 80% = 6720
Actually redeemed = 7500
Premium liability = [(2800 + 6720 - 7500) x ($2.50 + $7.50 - $6.00)] = 2020 x $4 = $8080